# Diabatic error and propagation of Majorana zero modes in interacting quantum dots systems

### v1

@@ Abstract
Motivated by recent experimental progress in realizing Majorana zero modes (MZMs) using quantum dot systems, we investigate the diabatic errors associated with the movement of those MZMs. The movement is achieved by tuning time-dependent gate potentials applied to individual quantum dots, effectively creating a moving potential wall. To probe the optimized movement of MZMs, we calculate the experimentally accessible time-dependent fidelity and local density-of-states using many-body time-dependent numerical methods. Our analysis reveals that an optimal potential wall height is crucial to preserve the well-localized nature of the MZM during its movement. Moreover, for the first time, we analyze diabatic errors in realistic quantum-dot systems, incorporating the effects of repulsive Coulomb interactions and disorder in both hopping and pairing terms. Additionally, we provide a comparative study of diabatic errors arising from the simultaneous versus sequential tuning of multiple gates during the MZMs movement. Finally, we estimate the time scale required for MZM transfer in a six-quantum-dot system, demonstrating that MZM movement is feasible and can be completed well within the qubit’s operational lifetime in practical quantum-dot setups.
@@ I Introduction
Majorana zero modes (MZMs) have potential applications in the development of fault-tolerant quantum computation Kitaev1 ; Kitaev2 ; brien ; Sarma ; Nayak . The non-local encoding of information and non-Abelian exchange statistics of MZMs allow for the construction of topologically protected qubits and gates, which are immune to local perturbations and decoherence. Experimentally achieving Majorana zero modes has been studied extensively through a hybrid semiconductor nanowire or a ferro-magnetic atomic chain system with strong spin-orbit coupling placed over a superconductor Roman ; Mourik . However, in these systems the presence of disorder and impurities continues to pose significant challenges to the formation and detection of Majorana zero modes Stanescu ; Gomanko ; Loo ; PanH . Recently, a novel approach has been proposed to realize Majorana zero modes (MZMs) by utilizing a chain of quantum dots Jay ; Loss ; Deng . The experimental realization of the minimal Kitaev chain using just two quantum dots coupled with a short superconductor-semiconductor hybrid, opened a new platform to observe MZMs Dvir . Remarkably, using three quantum dots, and working at the “sweet spot” $t_{h}=\Delta$ (where the electronic hopping $t_{h}$ and superconducting coupling $\Delta$ are equal in magnitude) ChunX two peaks in tunneling conductance measurements were observed interpreted as localized MZMs. Interestingly, the topological protection of MZMs increased compared to the two-site Kitaev experiments Bordin ; ChunX . Very recently, an enhancement of the excitation gap using Yu-Shiba-Rusinov states was shown experimentally in quantum-dot systems Zatelli .
@@
The above described quantum-dot approach provides well-controlled setups where the effect of disorder Dvir ; Bordin ; ChunX ; Zatelli is significantly reduced compared to the nanowire systems. These quantum-dot systems offer a promising platform for quantum information processing with Majorana zero modes (MZMs) Liu ; Mazur , where the manipulation of MZMs can be precisely controlled using local gates. The optimal movement of Majorana zero modes (MZMs) is crucial for braiding or fusion of MZMs in the broad framework of topological quantum computation Alicea ; Aasen . However, an important issue is that the manipulation of MZMs in quantum-dot systems for such applications must be completed within the quasiparticle “poisoning” time of the MZMs ChunX . Specifically, a rapid movement of MZMs can induce quasiparticle excitations between the ground-state manifold and the excited states, resulting in the loss of quantum information. Therefore, it is essential to manipulate MZMs within a finite window in time that (i) makes quantum computation doable, while simultaneously (ii) minimizing unwanted diabatic transitions MENG ; perfetto ; LXU .
@@
Figure 1: Schematic representation of the movement of the right edge MZM $\gamma_{2}$ in a coupled quantum-dots system using individually tunable gates attached to each quantum dot. (a) At time $t=0$, the system has two MZMs ($\gamma_{1}$, $\gamma_{2}$) located on the left and right edges of the system at the sweet-spot $t_{h}=\Delta$. (b) At time $t=t_{0}$, the right edge MZMs moved to site $i=i_{0}$ by using a time-dependent onsite chemical potential $\mu(t)$ ($\mu_{0}$ is the height of the potential wall). The onsite chemical potential can be tuned by application of local gates attached to quantum-dots with quench rate $\tau$. The blue color regions are topological while the red color regions are non-topological
@@
To study diabatic errors during the movement of MZMs in the quantum-dot system, we calculate the time-dependent fidelity Truong as well as the time-dependent local density-of-states (LDOS) for each quantum dot Bpandey1 ; Bpandey2 . First, we focus on the height of the potential wall required to move the localized MZM from one quantum dot to another. We find that optimizing the wall height is crucial to maintaining the well-localized behavior of the MZM, and measuring the local density-of-states plays a vital role in this optimization. Second, we analyze the effects of the nearest-neighbor Coulomb interaction on diabatic errors during the movement of MZMs. In realistic quantum dot systems the hopping and pairing terms may not be perfectly equal for a moderate number of quantum dots, leading to disorder in these terms ChunX ; Jay . For this reason, we investigate the diabatic error in the presence of disorder in the hopping and pairing terms and find that the diabatic error increases very slowly for moderate disorder. However, a further increase in disorder strength leads to a jump in the diabatic error. Third, we investigate the diabatic error during MZM movement in quantum-dot systems for the case when multiple gates are tuned simultaneously. Our results show that for slower tuning rates of the gate potentials, the diabatic error is significantly reduced in the case of multiple-gate tuning compared to single-gate tuning. Finally, using six quantum-dots, we estimate the time scale to move the MZMs in realistic quantum-dot experimental setups ChunX .
The organization of the manuscript is as follows. Section II introduces the interacting Kitaev model for quantum-dot systems at the sweet spot and provides details about our numerical simulations. Section III presents the results of our numerical simulations for the movement of MZMs, including the effect of the wall height (Sec. III.1), nearest-neighbor Coulomb repulsion (Sec. III.2), and disorder in the hopping and pairing terms (Sec. III.3). Section III.4 compares the diabatic error during the movement of MZMs for simultaneous and sequential gate tuning. Finally, Section IV provides our conclusions and discussion. In the Appendices, we include results on the non-analytic behavior of the many-body gap (Appendix A) and the diabatic error during MZM movement in the presence of disorder in the on-site potential (Appendix B).
@@ II Model and Methods
We start with the one-dimensional (1D) Kitaev model for spinless fermions under open boundary condition (OBC)
@@
where $c_{i}(c_{i}^{\dagger})$ is the fermionic anhilation (creation) operator, $n_{i}=c_{i}^{\dagger}c_{i}$ is the number operator, and $t_{h}$ and $\Delta$ are the nearest-neighbour hopping and SC pairing amplitudes, respectively. Throughout the calculation, we choose $t_{h}=1$ and $\Delta=t_{h}$ (unless stated otherwise) which is a good approximation for quantum-dot systems Dvir . Using the Hamiltonian in Eq. 1, we calculate the ground state using exact diagonalization (ED) Bpandey1 and/or state-of-art density matrix renormalization group (DMRG) Alvarez1 ; Nocera . At time $t=0$, for the parameters $t_{h}=\Delta=1$, $\mu_{i}=0$, and $V=0$, we have two onsite localized MZMs ($\gamma_{1}$ and $\gamma_{2}$) on the left and right edges of the Kitaev chain (see Fig. 1(a)). The right edge MZM ($\gamma_{2}$) is moved using sequential operations of the time-dependent gate potential (Fig. 1(b)). We introduce the time dependence by adding a time-dependent chemical potential in the Hamiltonian as follows Bpandey1
@@
$\mu_{i}(t)$ $=$ $\begin{dcases}0,&i<i_{0}\\ \mu_{0}\frac{n\Delta t}{\tau},&\text{if }i=i_{0}\\ \mu_{0},&i>i_{0}\end{dcases}$ (3)
where $\Delta t=0.001$ is the small time step used, $1/\tau$ is the quenched rate, and $n$ is the integer number of those steps chosen such that the chemical potential at site $i=i_{0}$ changes almost linearly with time from 00 to $\mu$ in a time $\tau$, which is equivalent to the gate switching time in real experimental setups. The time-dependent onsite chemical potential can be visualized as a moving wall which grows on site $i=i_{0}$ in time $\tau$ and then moves to the next site and grows again, as shown in Fig. 1. This effectively results in the movement of the MZM on the right edge towards the center, while keeping the MZM at the left edge intact.
@@ III Results
We present the results for the movement of Majorana zero modes (MZMs) in quantum-dot systems, focusing on the effects of the wall height, the strength of the nearest-neighbor Coulomb interactions, and the effect of disorder in the hopping and pairing terms. Furthermore, we compare the diabatic error during the MZM movement when multiple gates are tuned simultaneously.
@@ III.1 Effect of wall height during the movement of MZMs
In this subsection, we analyze the effect of wall height on the movement of the right edge MZM ($\gamma_{2}$), by varying the constant term $\mu_{0}$ in Eq. 3. To move the MZM from one site to another within a finite time $t$, the onsite chemical potential, $\mu_{i}(t)$, must be larger than the critical value $\mu_{c}>2t_{h}$ Alicea , in order to switch a site from a topologically nontrivial phase to a trivial phase. To track the movement of the MZM, we calculate both the time-dependent fidelity and the local density-of-states. At $t=0$, we choose parameters $t_{h}=\Delta=1$, $L=12$, and $V=0.05$, resulting in two localized MZMs $(\gamma_{1},\gamma_{2})$ at the left and right edges of the chain, as shown in Fig. 2(a). The site-dependent local density-of-states, $LDOS(\omega,j,t)$, at $t=0$, exhibits sharp peaks at sites $j=1$ and $12$, indicating two localized MZMs on the left and right edges of the Kitaev chain (Fig. 2(b)). The inset of Fig. 2(b) shows the electron and hole part of the local density-of-states $LDOS(\omega)$ as a function of $\omega$ for the right edge site $j=12$ and at $t=0$. The equal-height sharp peaks of the electron and hole components of $LDOS(\omega)$ at $\omega=0$ confirm the presence of a localized $\gamma_{2}$, on the right edge of the chain  Bpandey3 ; Herbrych .
As described in the previous section, the right-edge Majorana zero mode (MZM), $\gamma_{2}$, can be moved via the sequential modulation of the onsite chemical potential, $\mu_{i}(t)$. In Fig. 2, we examine the diabatic error, $f(t)$, for three different cases where $\gamma_{2}$ is moved using wall heights with $\mu_{0}=3$, $6$, and $12$. The MZM at the right edge is transported from site $j=12$ to $j=6$ in the same total time $T$ for the three cases. As shown in Fig. 2(c), the diabatic error is the smallest for $\mu_{0}=3$ and increases with higher wall heights. This behavior arises because the construction of higher walls requires a faster rate of change in $\mu_{i}(t)$ compared to lower walls. The sudden change in chemical potential at the start of wall formation at each site $j$ induces oscillations in the diabatic error, $f(t)$. For faster changes in $\mu_{i}(t)$, these oscillations exhibit larger amplitudes due to the more pronounced nonanalytic behavior associated with a rapid modulation of the chemical potential Truong ; Garraway ; Vitanov . Furthermore, we observe that the effective gap (defined as the energy difference between the ground state and the first excited state) shows discontinuities only at the beginning of increasing the on-site chemical potential at each site (see Fig. 6 in the Appendix A).

### v2

@@ Abstract
Motivated by recent experimental progress in realizing Majorana zero modes (MZMs) using quantum dot systems, we investigate the diabatic errors associated with the movement of those MZMs. The movement is achieved by tuning time-dependent gate potentials applied to individual quantum dots, effectively creating a moving potential wall. In order to probe the optimized movement of MZMs, we calculate the experimentally accessible time-dependent fidelity and local density-of-states using many-body time-dependent numerical methods. Our analysis reveals that an optimal potential wall height is crucial to preserve the well-localized nature of the MZM during its movement. Moreover, for the first time, we analyze diabatic errors in realistic quantum-dot systems, incorporating the effects of repulsive Coulomb interactions and disorder in both hopping and pairing terms. Additionally, we provide a comparative study of diabatic errors arising from the simultaneous versus sequential tuning of multiple gates during the MZMs movement. Finally, we estimate the time scale required for MZM transfer in a six-quantum-dot system, demonstrating that MZM movement is feasible and can be completed well within the qubit’s operational lifetime in practical quantum-dot setups.
@@ I Introduction
Majorana zero modes (MZMs) have potential applications in the development of fault-tolerant quantum computation Kitaev1 ; Kitaev2 ; brien ; Sarma ; Nayak . The non-local encoding of information and non-Abelian exchange statistics of MZMs allow for the construction of topologically protected qubits and gates, which are immune to local perturbations and decoherence. Experimentally achieving Majorana zero modes has been studied extensively through a hybrid semiconductor nanowire or a ferro-magnetic atomic chain system with strong spin-orbit coupling placed over a superconductor Roman ; Mourik . However, in these systems the presence of disorder and impurities continues to pose significant challenges to the formation and detection of Majorana zero modes Stanescu ; Gomanko ; Loo ; PanH . Recently, a novel approach has been proposed to realize Majorana zero modes (MZMs) by utilizing a chain of quantum dots Jay ; Loss ; Deng . The experimental realization of the minimal Kitaev chain using just two quantum dots coupled via a short superconductor-semiconductor hybrid Dvir , has given new hope for the observation of Majorana zero modes. In these systems, the effective couplings between quantum dots are gate-tunable and mediated by well-controlled Andreev bound states in the hybrid semiconductor-superconductor region. The excellent local control over the chemical potential of each quantum dot significantly reduces the effects of disorder and smooth gate potentials compared to more conventional proximitized semiconductor nanowire systems  Dvir ; Bordin ; ChunX ; Zatelli .
The conventional nano-wire SC heterostructures left several challenges open, such as to pin down the correct effective low-energy model. Meanwhile, the microscopic modeling of the new quantum-dot platform is tentatively simpler and more closely tied to the original Kitaev toy model. Very recently, using three quantum dots, and working at the “sweet spot” $t_{h}=\Delta$ (where the electronic hopping $t_{h}$ and superconducting coupling $\Delta$ are equal in magnitude) ChunX two peaks in tunneling conductance measurements were observed interpreted as localized MZMs. The topological protection of MZM increased compared to the Kitaev experiments with two sites Bordin ; ChunX . In the quantum-dot system the MZMs are perfectly localized onsite at the “sweet spot” $t_{h}=\Delta$ Remarkably, an enhancement of the excitation gap using Yu-Shiba-Rusinov states was shown experimentally in quantum-dot systems Zatelli . These quantum-dot systems offer a promising platform for quantum information processing with Majorana zero modes (MZMs) Liu ; Mazur , where the manipulation of MZMs can be precisely controlled using local gates. The controlled movement of MZMs across three quantum dots has been experimentally realized recently by tuning the chemical potential of the outer dots Bas .
@@
The optimal transport of Majorana zero modes (MZMs) is crucial for braiding or fusion of MZMs in the broad framework of topological quantum computation Alicea ; Aasen ; Rachel1 . However, an important issue is that the manipulation of MZMs in quantum-dot systems for such applications must be completed within the quasiparticle “poisoning” time of the MZMs ChunX . Specifically, a rapid movement of MZMs can induce quasiparticle excitations between the ground-state manifold and the excited states, resulting in the loss of quantum information. Therefore, it is essential to manipulate MZMs within a finite window in time that (i) makes quantum computation doable, while simultaneously (ii) minimizing unwanted diabatic transitions MENG ; perfetto ; LXU .
@@
Figure 1: Schematic representation of the movement of the right edge MZM $\gamma_{2}$ in a coupled quantum-dots system using individually tunable gates attached to each quantum dot. (a) At time $t=0$, the system has two MZMs ($\gamma_{1}$, $\gamma_{2}$) located on the left and right edges of the system at the sweet-spot $t_{h}=\Delta$. (b) At time $t=t_{0}$, the right edge MZMs moved to site $i=i_{0}$ by using a time-dependent onsite chemical potential $\mu(t)$ (($\mu_{0}$ is the height of the potential wall)). The onsite chemical potential can be tuned by application of local gates attached to quantum-dots with quench rate $\tau$. The blue color regions are topological while the red color regions are non-topological
@@
To study diabatic errors during the movement of MZMs in the quantum-dot system, we calculate the time-dependent fidelity Truong as well as the time-dependent local density-of-states (LDOS) for each quantum dot Bpandey1 ; Bpandey2 . First, we focus on the height of the potential wall required to move the localized MZM from one quantum dot to another. We find that optimizing the wall height is crucial to maintaining the well-localized behavior of the MZM, and measuring the local density-of-states plays a vital role in this optimization. Second, we analyze the effects of the nearest-neighbor Coulomb interaction on diabatic errors during the movement of MZMs. In realistic quantum dot systems the hopping and pairing terms may not be perfectly equal for a moderate number of quantum dots, leading to disorder in these terms ChunX ; Jay . For this reason, we investigate the diabatic error in the presence of disorder in the hopping and pairing terms and find that the diabatic error increases very slowly for moderate disorder. However, a further increase in disorder strength leads to a jump in the diabatic error. Third, we investigate the diabatic error during MZM movement in quantum-dot systems for the case when multiple gates are tuned simultaneously, using both linear and sine-squared ramping protocols. We find that the sine-squared ramping protocol is much more efficient and produces significantly less error than the linear ramp protocol. Our results show that for slower tuning rates of the gate potentials, the diabatic error is significantly reduced in the case of multiple-gate tuning compared to single-gate tuning. Finally, using six quantum-dots, we estimate the time scale to move the MZMs in realistic quantum-dot experimental setups ChunX .
The organization of the manuscript is as follows. Section II introduces the interacting Kitaev model for quantum-dot systems at the sweet spot and provides details about our numerical simulations. Section III presents the results of our study for the movement of MZMs, including the effect of the wall height (Sec. III.1), nearest-neighbor Coulomb repulsion (Sec. III.2), and disorder in the hopping and pairing terms (Sec. III.3). Section III.4 compares the diabatic error during the movement of MZMs for simultaneous and sequential gate tuning. Finally, Section IV provides our conclusions and discussion. In the Appendices, we include results on the non-analytic behavior of the many-body gap (Appendix A) and the diabatic error during MZM movement in the presence of disorder in the on-site potential (Appendix B).
@@ II Model and Methods
The recent progresses in quantum-dot systems have enabled the experimental realization of two- and three-sites Kitaev chains Dvir ; ChunX . In these setups, spin-polarized quantum dots and short superconductor-semiconductor hybrids are coupled by elastic co-tunnelling (ECT) and crossed Andreev reflection (CAR) Dvir ; ChunX . This configuration allows for tunable single-electron hopping $t$ (ECT) and triplet pairing $\Delta$ (CAR), both mediated by Andreev bound states in the hybrid segments Liu . These couplings in the quantum-dots systems are tuned by electrostatic gates connected to the SC-SM hybrid segment and provides all the key ingredients for implementing artificial Kitaev chains Dvir . The Kitaev chain Hamiltonian for spinless fermions, with hopping $t$ and $p$-wave pairing $\Delta$, can be described as Kitaev1 :
@@
where $c_{i}(c_{i}^{\dagger})$ is the fermionic anhilation (creation) operator, $n_{i}=c_{i}^{\dagger}c_{i}$ is the number operator, and $t_{h}$ and $\Delta$ are the nearest-neighbour hopping and SC pairing amplitudes, respectively. Throughout the calculation, we set $t_{h}=1$ and $\Delta=t_{h}$ (unless stated otherwise), a condition that has been realized in quantum-dot systems Dvir . In the case of a two-site quantum-dot system, two perfectly localized MZMs were observed at the sweet spot $2t_{h}=2\Delta=25\,\mu eV$. The tuning of the coupling ratio $t_{h}/\Delta=1$ is easily achieved by adjusting the electrostatic gate potential in the quantum-dot system, which enforces the product-state property of the ground state and thereby yields maximally localized Majorana states Thomale2 . Using the Hamiltonian in Eq. 1, we calculate the ground state using exact diagonalization (ED) Bpandey1 and/or state-of-art density matrix renormalization group (DMRG) Alvarez1 ; Nocera . At time $t=0$, for the parameters $t_{h}=\Delta=1$, and $V=0$, we have two onsite localized MZMs ($\gamma_{1}$ and $\gamma_{2}$) on the left and right edges of the Kitaev chain (see Fig. 1(a)). The right edge MZM ($\gamma_{2}$) is moved using sequential operations of the time-dependent gate potential (Fig. 1(b)). We introduce the time dependence by adding a time-dependent chemical potential in the Hamiltonian as follows Bpandey1
@@
$\mu_{i}(t)$ $=$ $\begin{dcases}0,&i<i_{0}\\ \mu_{0}\frac{t}{\tau},&\text{if }i=i_{0}\\ \mu_{0},&i>i_{0}\end{dcases}$ (3)
where $1/\tau$ is the quench rate. We discretize the time steps as $t=k\Delta t$, where $\Delta t=0.001$ is a small time step, and $k$ is an integer chosen such that the chemical potential at site $i=i_{0}$ varies nearly linearly from 00 to $\mu_{0}$ over a time $\tau$. We first choose a linear time-dependent chemical potential since, in quantum-dot experiments for the two-site Kitaev model, the chemical potentials of quantum dots are linearly tuned by voltages applied via gates Dvir . We also compare the sine-squared ramping protocol with the linear ramping protocol. The time-dependent onsite chemical potential can be visualized as a moving potential wall that grows at site $i=i_{0}$ over time $\tau$, then shifts to the next site $i=i_{0}-1$, and grows again, as illustrated in Fig. 2. This process is analogous to gate switching in real experimental setups. As a result, the MZM on the right edge moves toward the center, while the MZM on the left edge remains localized (Fig. 1).
@@
Figure 2: Ramping protocol to move a MZM through the system. At time $t/\tau=0$, we prepare the system in the ground state with all the onsite potential set at 00. For a time interval $0\leq t/\tau\leq 1$, the onsite potential at the right most site (site $i=12$) is increased from 00 to $\mu_{0}$ linearly and stays constant on-wards. Later on, in the time interval $1\leq t/\tau\leq 2$, the onsite potential at the rightmost site remains constant, but the onsite potential at the next site ($i=11$) increases linearly until it reaches $\mu_{0}$.
@@
In addition to fidelity, we also calculate the time-dependent local density-of-states (at zero temperature). As described in Ref. Bpandey1 , first we time evolve the initial state $|\psi(0)\rangle$ up to time $t=N_{R}\tau$, (here $N_{R}$ is the number of sites where the chemical potential reaches its maximum value $\mu_{0}$) using the time-dependent Hamiltonian $H(t)$ as: $|\Psi(t)\rangle=U(t)|\psi(0)\rangle$. In order to obtain the time and energy resolved local density-of-states, we calculate the double-time Green function $G(t,t^{\prime})$ Kennes , using the instantaneous Hamiltonian $H_{f}=H(t=t_{f})$ at time $t_{f}=N_{R}\tau$:
@@ III Results
In this section, we present the results for the movement of Majorana zero modes (MZMs) in quantum-dot systems, focusing on the effects of the wall height, the strength of the nearest-neighbor Coulomb interactions, and the effect of disorder in the hopping and pairing terms. Furthermore, we compare the diabatic error during the MZM movement when multiple gates are tuned simultaneously.
@@ III.1 Effect of wall height during the movement of MZMs
In this subsection, we analyze the effect of wall height on the movement of the right edge MZM ($\gamma_{2}$), by varying the constant term $\mu_{0}$ in Eq. 3. To move the MZM from one site to another within a finite time $t$, the onsite chemical potential, $\mu_{i}(t)$, must be larger than the critical value $\mu_{c}>2t_{h}$ Alicea , in order to switch a site from a topologically nontrivial phase to a trivial phase. To track the movement of the MZM, we calculate both the time-dependent fidelity and the local density of states. At $t=0$, we choose parameters $t_{h}=\Delta=1$, $L=12$, and $V=0.05$, resulting in two localized MZMs $(\gamma_{1},\gamma_{2})$ at the left and right edges of the chain, as shown in Fig. 3(a). The site-dependent local density-of-states, $LDOS(\omega,j,t)$, at $t=0$, exhibits sharp peaks at sites $j=1$ and $12$, indicating two localized MZMs on the left and right edges of the Kitaev chain (Fig. 3(b)). The inset of Fig. 3(b) shows the electron and hole part of the local density-of-states $LDOS(\omega)$ as a function of $\omega$ for the right edge site $j=12$ and at $t=0$. The equal-height sharp peaks of the electron and hole components of $LDOS(\omega)$ at $\omega=0$ confirm the presence of a localized $\gamma_{2}$, on the right edge of the chain  Bpandey3 ; Bpandey4 ; Herbrych .
As described in the previous section, the right-edge Majorana zero mode (MZM), $\gamma_{2}$, can be moved via the sequential modulation of the onsite chemical potential, $\mu_{i}(t)$. In Fig. 3, we examine the diabatic error, $f(t)$, for three different cases where $\gamma_{2}$ is moved using wall heights with $\mu_{0}=3$, $6$, and $12$. The MZM at the right edge is transported from site $j=12$ to $j=6$ in the same total time $T$ for the three cases. As shown in Fig. 3(c), the diabatic error is the smallest for $\mu_{0}=3$ and increases with higher wall heights. This behavior arises because the construction of higher walls requires a faster rate of change in $\mu_{i}(t)$ compared to lower walls (for the cases when the total ramp time $\tau$ is kept fixed). The sudden change in chemical potential at the start of wall formation at each site $j$ induces oscillations in the diabatic error, $f(t)$. For faster changes in $\mu_{i}(t)$, these oscillations exhibit larger amplitudes due to the more pronounced nonanalytic behavior associated with a rapid modulation of the chemical potential Truong ; Garraway ; Vitanov . Furthermore, we observe that the effective gap (defined as the energy difference between the ground state and the first excited state) shows discontinuities only at the beginning of increasing the on-site chemical potential at each site (see Fig. 7 in the Appendix A).